# FLRW embeddings in $\mathbb{R}^{n+2}$, differential geometry and conformal photon propagator

### v1

@@ I Introduction
Regarding the relationship between the intrinsic and ambient quantities we note that it is fundamentally a matter of differential geometry. For instance, one has to relate operators such as the Laplace-de Rham differential operator, viewed as the wave operator for the physical field, to its ambient realization. There is then an interplay to analyze between restriction to and extension out of the embedded submanifold of differential operators. Such a program has already been carried out successfully for scalar fields and one-forms,[10, 11, 12, 13] in particular to (A)dS spaces. The current work encompasses these previous results, while vastly extending and simplifying them. In our framework, the spacetime $X_{f}$ is obtained as the intersection of the null cone $\mathcal{C}$ of $\mathbb{R}^{n+2}$ with an hypersurface $\mathcal{P}_{f}$, defined by a homogeneous function $f$ of degree one in ambient space. We show, with explicit formulas, that every differential-geometric objects needed such as the curvature tensors, the Laplace-de Rham $\square_{f}$, and related entities, are derived solely from this defining function $f$.
@@ II.2.2 The general case for $X_{f}$
with $T_{y}(\mathbb{R}^{n+2})\simeq\mathbb{R}^{n+2}$. First note that, at $y\in X_{f}$, $F$ and $D$ are two linearly independent vectors since $\eta(F,D)=\langle df,D\rangle=D(f)=f=1$ while $\eta(D,D)=2c(y)=0$, thus they span a two dimensional vector space. Then, vectors spanned by $F$ and $D$ are orthogonal, with respect to $\eta$, to $T_{y}(X_{f})$ by construction. Now, consider the vector $V=\alpha F+\beta D\in(T_{y}(X_{f})\cap\mathrm{span}(F,D))$, with $\alpha$ and $\beta$ two reals, then since $\eta(V,D)=\beta=0$ and $\eta(V,F)=\alpha=0$ one concludes that the intersection is reduced to the null vector. This establishes Eq. (1).
As a consequence there are no isotropic vectors in $T_{y}(X_{f})$ which are orthogonal to $T_{y}(X_{f})$ and $\eta_{f}(u,v)=m_{f}{}^{*}\eta(u,v)$, for $u,v\in T_{y}(X_{f})$ is non-degenerate. ∎
@@ III.4 The special case of FLRW spaces
Within $\mathbb{R}^{n+2}$, and according to Eq. (3), one can obtain each of the FLRW space by using the defining function $f_{k}$, not to be confused with $f_{\kappa}$ appearing in Eq. (3) defining AdSM spaces, as
@@ IV.1 A convenient basis of $T_{y}(\mathbb{R}^{n+2})$ around $X_{f}$
$e^{n+1}=\eta^{n+1\,n+1}\flat_{\eta}e_{n+1}=df+\frac{1}{2}(1-\mathbf{F}^{2})dc.$
@@ IV.2 Transverse and longitudinal ambient differential forms
$T:=i_{\scriptscriptstyle{\,}\!F}i_{\scriptscriptstyle{\,}\!D}j^{df}j^{dc},$
then a differential form $\alpha\in\Omega^{a}(\mathbb{R}^{n+2})$ is a transverse differential form with respect to $X_{f}$ iff $T\alpha=\alpha$ or, equivalently, iff $i_{\scriptscriptstyle{\,}\!F}\alpha=0$ and $i_{\scriptscriptstyle{\,}\!D}\alpha=0$. The set of transverse forms is noted $\Omega_{t}(\mathbb{R}^{n+2})$.
@@
$m_{f}{}^{*}(\alpha)=\alpha_{|X_{f}}\mbox{ iff }\alpha\in\Omega_{t}(\mathbb{R}^{n+2}).$
@@
A differential form $\alpha\in\Omega^{a}(\mathbb{R}^{n+2})$ is said to be longitudinal with respect to $X_{f}$ iff $L\alpha=\alpha$. The set of longitudinal forms is noted $\Omega_{l}(\mathbb{R}^{n+2})$.
@@ IV.3 Strongly transverse ambient differential forms
The mapping $m_{f}{}^{*}:\ \Omega_{t}(\mathbb{R}^{n+2})\to\Omega(X_{f})$ is a surjection, we seek a smaller initial set to make it a bijection. Within the span of $F$ and $D$ we introduce a new vector field tangent to $\mathcal{P}_{f}$.
@@
Let $\beta$ be a differential form on $X_{f}$, and $r\in\mathbb{R}$, then at least locally, there exists a unique differential form $\hat{\beta}$ extending $\beta$ (i.e. $m_{f}{}^{*}(\hat{\beta})=\beta)$ in a neighborhood of $X_{f}\subset\mathbb{R}^{n+2}$ such that $(\mathscr{L}_{\scriptscriptstyle{\!}D}-r\mathrm{{Id}})\hat{\beta}=\mathscr{L}_{\scriptscriptstyle{E}_{f}}\hat{\beta}=0$. Moreover, $\hat{\beta}$ is transverse (i.e. $i_{\scriptscriptstyle{\,}\!D}\hat{\beta}=i_{\scriptscriptstyle{\,}\!F}\hat{\beta}=0$).
@@
The condition $(\mathscr{L}_{\scriptscriptstyle{\!}D}-r\mathrm{{Id}})\hat{\beta}=0$ is equivalent to $\hat{\beta}$ is homogeneous of degree $r$.
@@ V.1 Pull-back of differential operators on $p$-forms
$m_{f}{}^{*}(d\alpha)=d_{f}m_{f}{}^{*}(\alpha)=d_{f}\alpha_{f},$ (17)
@@ VII.1 The massless conformally coupled scalar field
To begin in the simplest case conceivable, let us consider the equation of a massless scalar field coupled to the scalar curvature fulfilling, using Eqs. (19) and (25), we obtain
@@
with $y=m_{f}(x)$, $y^{\prime}=m_{f}(x^{\prime})$ and $y\cdot y^{\prime}$ the ambient product between two points in $\mathbb{R}^{6}$. We denote the two points in $X_{f}$ as $x$ and $x^{\prime}$ in the two-point function to maintain consistency with standard notations. It is to be understood that some $i\epsilon$ must be added to regularize the function. Exploiting the homogeneity of the field (of degree 1-1), this two-point function can then be rescaled from $X_{f}$ to the null cone $\mathcal{C}$. In this way, one recognizes that the ambient expression (32) of the two-point function is valid for arbitrary conformally flat spacetimes $X_{f}$.
@@ A.1 Summary of our notations in differential geometry
Let $(M,g)$ be a $n$-dimensional oriented pseudo-Riemannian manifold with $\nabla^{g}$ its Levi-Civita connection. We note $\Omega^{p}(M)=\smash{\bigwedge^{p}T^{*}M}$ and $\Omega(M)=\sum_{p}\Omega^{p}(M)$ the sets of differential forms on $M$ and $\tilde{g}$ the corresponding metric on $\Omega(M)$. The volume form is written $\omega_{g}$ and the Hodge operator $_{g}$ is defined, as usual, through $\alpha_{1}\wedge*_{g}\alpha_{2}=\tilde{g}(\alpha_{1},\alpha_{2})\omega_{g}$ where $\alpha_{1},\alpha_{2}\in\Omega^{p}(M)$. The natural pairing between a $1$-form $\lambda$ and a vector $v$ is written $\langle\lambda,v\rangle$, the musical applications $\flat_{g}$ and $\sharp_{g}$ relate this pairing to the metric as $\langle\lambda,v\rangle=g(\sharp_{g}\lambda,v)$ and $g(u,v)=\langle\flat_{g}u,v\rangle.$
@@ A.2 Schouten-Nijenhuis bracket on forms
Let $\phi$ a scalar field, key formulas for the present article are
@@ B.3 Minkowski and AdS as alternative base spaces for k=1k=-1
By scaling the coordinates described in Eq. (71) with the factor $\Omega$, we retrieve the FLRW embedding of the k=1k=-1 case. This serves as a third method, via AdS, in obtaining the k=1k=-1 embedding given in Eq. (7).
@@ C.1 FLRW spaces with k=1k=-1
For the k=1k=-1 FLRW case whose defining function is given in Eq. (9) one has
@@
with $y^{n+1}/y^{n}=\tanh(t)$. Then, $J(f_{k=-1})=0$ splits into three cases: (i) $J^{0n}=0$ and $J^{0n+1}\neq 0$, (ii) $J^{0n}\neq 0$ and $J^{0n+1}=0$, (iii) $J^{0n}\neq 0$ and $J^{0n+1}\neq 0$. In the first case, one reaches an isometry if $\overset{\,{}_{\mbox{\Large.}}}{a}+\tanh(t)a=0$ that is for $a(t)=1/\cosh(t)=\mathrm{sech}(t)$ thus recovering an AdS space [Eq. (71)] where the additional isometries are $y_{\mu}\partial_{n+1}-y_{n+1}\partial_{\mu}$ completing SO(1,n1)SO(1,n-1) to SO(2,n1)SO(2,n-1). The second case, is very similar to the first case in which one obtains additional isometries if $\overset{\,{}_{\mbox{\Large.}}}{a}+a/\tanh(t)=0$ that is for $a(t)=1/\sinh(t)=\mathrm{csch}(t)$ thus recovering a dS space [Eq. (66)] where the additional isometries are $y_{\mu}\partial_{n}-y_{n}\partial_{\mu}$ completing SO(1,n1)SO(1,n-1) to $SO(1,n)$. For the final case, let us set $q=J^{0n+1}/J^{0n}$ a non-zero real constant, then reaching an additional isometry amounts to solve
@@ C.2 FLRW spaces with $k=0$
For the $k=0$ FLRW case whose defining function is given in Eq. (9) one has
@@
with $t=y^{0}/(y^{n}+y^{n+1})$. The equation $J(f)=0$ is obviously fulfilled by the $\tfrac{1}{2}(n-1)(n-2)$ rotations generated by $y_{i}\partial_{j}-y_{j}\partial_{i}$ and by the (n1)(n-1) translations $P_{i}=\tfrac{1}{2}[y_{i}(\partial_{n}-\partial_{n+1})-(y_{n}-y_{n+1})\partial_{i}]$, together they generate the (minimal) group of isometries E(n1)E(n-1) of the $k=0$ FLRW spaces. Finding all of the solutions to the above equation is again rather straightforward. First by considering the case of $\alpha=i$ one has three cases to consider: (i) $J^{0i}=0$ and $(J^{in}+J^{in+1})\neq 0$, (ii) $J^{0i}\neq 0$ and $(J^{in}+J^{in+1})=0$, (iii) $J^{0i}\neq 0$ and $(J^{in}+J^{in+1})\neq 0$. In the first case, one has additional isometries iff $a$ is a solution to $t\overset{\,{}_{\mbox{\Large.}}}{a}+a=0$, that is for $a(t)=t^{-1}$ which corresponds to a dS space [Eq. (67)]. To the generators of E(n1)E(n-1) one can then add the (n1)(n-1) extra generators $T_{i}=\tfrac{1}{2}[y_{i}(\partial_{n}+\partial_{n+1})-(y_{n}+y_{n+1})\partial_{i}]$ and the hyperbolic rotation $y_{n}\partial_{n+1}-y_{n+1}\partial_{n}$. Thus, considering then $P_{i}\pm T_{i}$, we recover the full de Sitter group $SO(1,n)$ with the generators $y_{i}\partial_{j}-y_{j}\partial_{i}$, $y_{n}\partial_{i}-y_{i}\partial_{n}$, $y_{n+1}\partial_{i}-y_{i}\partial_{n+1}$ and $y_{n}\partial_{n+1}-y_{n+1}\partial_{n}$, that is the usual ones in which the roles of $y^{0}$ and $y^{n+1}$ have been swapped. In the second case, one has additional isometries if $\overset{\,{}_{\mbox{\Large.}}}{a}=0$, that is $a=1$ and the FLRW space simply is Minkowski space [Eq. (62)]. Then, to E(n1)E(n-1) one can add the n1n-1 hyperbolic rotations $y_{0}\partial_{i}-y_{i}\partial_{0}$ and the translation $P_{0}=\frac{1}{2}[y_{0}(\partial_{n}-\partial_{n+1})-(y_{n}-y_{n+1})\partial_{0}]$ thus obtaining the Poincaré group. For the final case, we set $q=J^{0i}/(J^{in}+J^{in+1})$ and the equation to solve to obtain additional isometries is $(t+q)\overset{\,{}_{\mbox{\Large.}}}{a}(t)+a(t)=0$, whose solution is another dS space, which we studied in case (i). Now, for $\alpha\neq i$, one has to find the most general solutions to the system

### v2

@@ I Introduction
Regarding the relationship between intrinsic and ambient quantities we note that it is fundamentally a matter of differential geometry. For instance, one has to relate operators such as the Laplace-de Rham differential operator, viewed as the wave operator for the physical field, to its ambient realization. There is then an interplay to analyze between restriction to and extension out of the embedded submanifold of differential operators. Such a program has already been carried out successfully for scalar fields and one-forms,[10, 11, 12, 13] in particular to (A)dS spaces. The current work encompasses these previous results, while vastly extending and simplifying them. In our framework, the spacetime $X_{f}$ is obtained as the intersection of the null cone $\mathcal{C}$ of $\mathbb{R}^{n+2}$ with an hypersurface $\mathcal{P}_{f}$, defined by a homogeneous function $f$ of degree one in ambient space. We show, with explicit formulas, that every differential-geometric objects needed such as the curvature tensors, the Laplace-de Rham $\square_{f}$, and related entities, are derived solely from this defining function $f$.
@@ II.2.2 The general case for $X_{f}$
with $T_{y}(\mathbb{R}^{n+2})\simeq\mathbb{R}^{n+2}$. First note that, at $y\in X_{f}$, $F$ and $D$ are two linearly independent vectors since $\eta(F,D)=\langle df,D\rangle=D(f)=f=1$ while $\eta(D,D)=2c(y)=0$, thus they span a two dimensional vector space. Then, vectors spanned by $F$ and $D$ are orthogonal to $T_{y}(X_{f})$, with respect to $\eta$, by construction. Now, consider the vector $V=\alpha F+\beta D\in(T_{y}(X_{f})\cap\mathrm{span}(F,D))$, with $\alpha$ and $\beta$ two reals, then since $\eta(V,D)=\beta=0$ and $\eta(V,F)=\alpha=0$ one concludes that the intersection is reduced to the null vector. This establishes Eq. (1).
As a consequence there are no isotropic vectors in $T_{y}(X_{f})$ which are orthogonal to $T_{y}(X_{f})$ and $\eta_{f}(u,v)=m_{f}{}^{*}\eta(u,v)$, for $u,v\in T_{y}(X_{f})$, is non-degenerate. ∎
@@ III.4 The special case of FLRW spaces
Within $\mathbb{R}^{n+2}$, and according to Eq. (3), one can obtain each type of FLRW space by using the defining function $f_{k}$, not to be confused with $f_{\kappa}$ appearing in Eq. (3) defining AdSM spaces, as
@@ IV.1 A convenient basis of $T_{y}(\mathbb{R}^{n+2})$ around $X_{f}$
$e^{n+1}=\eta^{n+1\,n+1}\flat_{\eta}e_{n+1}=df+\tfrac{1}{2}(1-\mathbf{F}^{2})dc.$
@@ IV.2 Transverse and longitudinal ambient differential forms
$T=i_{\scriptscriptstyle{\,}\!F}i_{\scriptscriptstyle{\,}\!D}j^{df}j^{dc},$
then a differential form $\alpha\in\Omega^{a}(\mathbb{R}^{n+2})$ is a transverse differential form with respect to $X_{f}$ iff $T\alpha=\alpha$ or, equivalently, iff $i_{\scriptscriptstyle{\,}\!F}\alpha=0$ and $i_{\scriptscriptstyle{\,}\!D}\alpha=0$. The set of transverse forms is noted $\Omega_{\rm t}(\mathbb{R}^{n+2})$.
@@
$m_{f}{}^{*}\alpha=\alpha_{|X_{f}}\mbox{ iff }\alpha\in\Omega_{\rm t}(\mathbb{R}^{n+2}).$
@@
A differential form $\alpha\in\Omega^{a}(\mathbb{R}^{n+2})$ is said to be longitudinal with respect to $X_{f}$ iff $L\alpha=\alpha$. The set of longitudinal forms is noted $\Omega_{\rm l}(\mathbb{R}^{n+2})$.
@@ IV.3 Strongly transverse ambient differential forms
The mapping $m_{f}{}^{*}:\ \Omega_{\rm t}(\mathbb{R}^{n+2})\to\Omega(X_{f})$ is a surjection, we seek a smaller initial set to make it a bijection. Within the span of $F$ and $D$ we introduce a new vector field tangent to $\mathcal{P}_{f}$.
@@
Let $\beta$ be a differential form on $X_{f}$, and $r\in\mathbb{R}$, then at least locally, there exists a unique differential form $\hat{\beta}$ extending $\beta$ (i.e. $m_{f}{}^{*}\hat{\beta}=\beta)$ in a neighborhood of $X_{f}\subset\mathbb{R}^{n+2}$ such that $(\mathscr{L}_{\scriptscriptstyle{\!}D}-r\mathrm{{Id}})\hat{\beta}=\mathscr{L}_{\scriptscriptstyle{E}_{f}}\hat{\beta}=0$. Moreover, $\hat{\beta}$ is transverse (i.e. $i_{\scriptscriptstyle{\,}\!D}\hat{\beta}=i_{\scriptscriptstyle{\,}\!F}\hat{\beta}=0$).
@@
The condition $(\mathscr{L}_{\scriptscriptstyle{\!}D}-r\mathrm{{Id}})\hat{\beta}=0$ is equivalent to $\hat{\beta}$ being homogeneous of degree $r$.
@@ V.1 Pull-back of differential operators on $p$-forms
$m_{f}{}^{*}(d\alpha)=d_{f}m_{f}{}^{*}\alpha=d_{f}\alpha_{f},$ (17)
@@ VII.1 The massless conformally coupled scalar field
To begin in the simplest case conceivable, let us consider the equation of a massless scalar field coupled to the scalar curvature, using Eqs. (19) and (25), we obtain
@@
with $y=m_{f}(x)$, $y^{\prime}=m_{f}(x^{\prime})$ and $y\cdot y^{\prime}$ the ambient product between two points in $\mathbb{R}^{6}$. We denote the two points in $X_{f}$ as $x$ and $x^{\prime}$ in the two-point function to maintain consistency with standard notations. Additionally, it is to be understood that some $i\epsilon$ must be added to regularize the function. Then, exploiting the homogeneity of the field (of degree 1-1), this two-point function can be rescaled from $X_{f}$ to the null cone $\mathcal{C}$. In this way, one recognizes that the ambient expression (32) of the two-point function is valid for arbitrary conformally flat spacetimes $X_{f}$.
@@ A.1 Summary of our notations in differential geometry
Let $(M,g)$ be a $n$-dimensional oriented pseudo-Riemannian manifold with $\nabla^{g}$ its Levi-Civita connection. We note $\Omega^{p}(M)=\smash{\bigwedge^{p}T^{*}(M)}$ and $\Omega(M)=\sum_{p}\Omega^{p}(M)$ the sets of differential forms on $M$ and $\tilde{g}$ the corresponding metric on $\Omega(M)$. The volume form is written $\omega_{g}$ and the Hodge operator $_{g}$ is defined, as usual, through $\alpha_{1}\wedge*_{g}\alpha_{2}=\tilde{g}(\alpha_{1},\alpha_{2})\omega_{g}$ where $\alpha_{1},\alpha_{2}\in\Omega^{p}(M)$. The natural pairing between a $1$-form $\lambda$ and a vector $v$ is written $\langle\lambda,v\rangle$, the musical applications $\flat_{g}$ and $\sharp_{g}$ relate this pairing to the metric as $\langle\lambda,v\rangle=g(\sharp_{g}\lambda,v)$ and $g(u,v)=\langle\flat_{g}u,v\rangle.$
@@ A.2 Schouten-Nijenhuis bracket on forms
Let $\phi$ be a scalar field, key formulas for the present article are
@@ B.3 Minkowski and AdS as alternative base spaces for k=1k=-1
By scaling the coordinates described in Eq. (71) with the factor $\Omega(t)$, we retrieve the FLRW embedding of the k=1k=-1 case. This serves as a third method, via AdS, in obtaining the k=1k=-1 embedding given in Eq. (7).
@@ C.1 FLRW spaces with k=1k=-1
For the k=1k=-1 case, whose defining function is given in Eq. (9), one has
@@
with $y^{n+1}/y^{n}=\tanh(t)$. Then, $J(f_{k=-1})=0$ splits into three cases: (i) $J^{0n}=0$ and $J^{0n+1}\neq 0$, (ii) $J^{0n}\neq 0$ and $J^{0n+1}=0$, (iii) $J^{0n}\neq 0$ and $J^{0n+1}\neq 0$. In the first case, one reaches an isometry if $\overset{\,{}_{\mbox{\Large.}}}{a}+\tanh(t)a=0$ that is for $a(t)=1/\cosh(t)=\mathrm{sech}(t)$ thus recovering an AdS space [Eq. (71)] where the additional isometries are $y_{\mu}\partial_{n+1}-y_{n+1}\partial_{\mu}$ completing SO(1,n1)SO(1,n-1) to SO(2,n1)SO(2,n-1). The second case is very similar to the first case, one obtains additional isometries if $\overset{\,{}_{\mbox{\Large.}}}{a}+a/\tanh(t)=0$ that is for $a(t)=1/\sinh(t)=\mathrm{csch}(t)$ thus recovering a dS space [Eq. (66)] where the additional isometries are $y_{\mu}\partial_{n}-y_{n}\partial_{\mu}$ completing SO(1,n1)SO(1,n-1) to $SO(1,n)$. For the final case, let us set $q=J^{0n+1}/J^{0n}$ a non-zero real constant, then reaching an additional isometry amounts to solve
@@ C.2 FLRW spaces with $k=0$
For the $k=0$ case, whose defining function is given in Eq. (9), one has
@@
with $t=y^{0}/(y^{n}+y^{n+1})$. The equation $J(f)=0$ is obviously fulfilled by the $\tfrac{1}{2}(n-1)(n-2)$ rotations generated by $y_{i}\partial_{j}-y_{j}\partial_{i}$ and by the (n1)(n-1) translations $P_{i}=\tfrac{1}{2}[y_{i}(\partial_{n}-\partial_{n+1})-(y_{n}-y_{n+1})\partial_{i}]$, together they generate the (minimal) group of isometries E(n1)E(n-1) of the $k=0$ FLRW spaces. Finding all of the solutions to the above equation is again rather straightforward. First by considering the case of $\alpha=i$ one has three cases to consider: (i) $J^{0i}=0$ and $(J^{in}+J^{in+1})\neq 0$, (ii) $J^{0i}\neq 0$ and $(J^{in}+J^{in+1})=0$, (iii) $J^{0i}\neq 0$ and $(J^{in}+J^{in+1})\neq 0$. In the first case, one has additional isometries iff $a$ is a solution to $t\overset{\,{}_{\mbox{\Large.}}}{a}+a=0$, that is for $a(t)=t^{-1}$ which corresponds to a dS space [Eq. (67)]. To the generators of E(n1)E(n-1) one can then add the (n1)(n-1) extra generators $T_{i}=\tfrac{1}{2}[y_{i}(\partial_{n}+\partial_{n+1})-(y_{n}+y_{n+1})\partial_{i}]$ and the hyperbolic rotation $y_{n}\partial_{n+1}-y_{n+1}\partial_{n}$. Thus, considering then $P_{i}\pm T_{i}$, we recover the full de Sitter group $SO(1,n)$ with the generators $y_{i}\partial_{j}-y_{j}\partial_{i}$, $y_{n}\partial_{i}-y_{i}\partial_{n}$, $y_{n+1}\partial_{i}-y_{i}\partial_{n+1}$ and $y_{n}\partial_{n+1}-y_{n+1}\partial_{n}$, that is the usual ones in which the roles of $y^{0}$ and $y^{n+1}$ have been swapped. In the second case, one has additional isometries if $\overset{\,{}_{\mbox{\Large.}}}{a}=0$, that is $a=1$ and the FLRW space simply is Minkowski space [Eq. (62)]. Then, to E(n1)E(n-1) one can add the n1n-1 hyperbolic rotations $y_{0}\partial_{i}-y_{i}\partial_{0}$ and the translation $P_{0}=\frac{1}{2}[y_{0}(\partial_{n}-\partial_{n+1})-(y_{n}-y_{n+1})\partial_{0}]$ thus obtaining the Poincaré group. For the final case, we set $t_{0}=-J^{0i}/(J^{in}+J^{in+1})$ and the equation to solve to obtain additional isometries is $(t-t_{0})\overset{\,{}_{\mbox{\Large.}}}{a}(t)+a(t)=0$, whose solution is another dS space, which we studied in case (i). Now, for $\alpha\neq i$, one has to find the most general solutions to the system